OpenStudy (some.random.cool.kid):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function?

OpenStudy (some.random.cool.kid):

Part B: What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (some.random.cool.kid):

i have no idea *shrugs*

OpenStudy (perl):

is f(n) = 8 ( 1.05)^n

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

does anyone understand it?

OpenStudy (misty1212):

HI!

OpenStudy (some.random.cool.kid):

hello

OpenStudy (misty1212):

\[f(n) = 8(1.05)^n \] right ?

OpenStudy (some.random.cool.kid):

yes

OpenStudy (perl):

the y intercept is the height when n= 0 , or at time = zero

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

thats part b though

OpenStudy (misty1212):

lol the plant grows exponentially!! 5% a day

OpenStudy (misty1212):

it starts at \(f(0)=8\) mm tall

OpenStudy (some.random.cool.kid):

5%? are you sure

OpenStudy (misty1212):

pretty sure dear, yes

OpenStudy (some.random.cool.kid):

5% is part A but what about part b

OpenStudy (misty1212):

i guess you are supposed to solve \[8(1.05)^n=11.26\] oh my dear you are confused

OpenStudy (misty1212):

i answered B and @perl answered B already

OpenStudy (some.random.cool.kid):

oh what was b i didnt see it

OpenStudy (misty1212):

and the answer to A is not 5% i just said it grew at 5% a day is all

OpenStudy (some.random.cool.kid):

What is a reasonable domain to plot the growth function?

OpenStudy (some.random.cool.kid):

i dont need you to solve

OpenStudy (some.random.cool.kid):

i need that

OpenStudy (some.random.cool.kid):

@Hero

OpenStudy (perl):

The domain are the set of inputs , here days. The study begins at n=0, and ends at what day ?

OpenStudy (some.random.cool.kid):

it was n number of days? so i dont know really

OpenStudy (perl):

the domain will be n between 0 and some number. We don't want n to be too big because that will imply that the plant grows very high due to exponential growth. And notice that in the study the plant only grew to 11.26 cm

OpenStudy (some.random.cool.kid):

yeah

OpenStudy (some.random.cool.kid):

i could divide right?

OpenStudy (perl):

we can solve this equation: 8(1.05)^n=11.26 to give us a rough upper bound for n

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

8 x 1.05^2 =n?

OpenStudy (some.random.cool.kid):

then divide 11.26 by the number?

OpenStudy (perl):

not quite

OpenStudy (some.random.cool.kid):

oh

OpenStudy (perl):

8(1.05)^n=11.26 first divide both sides by 8 (1.05)^n = 11.26/8

OpenStudy (some.random.cool.kid):

oh i was close :P

OpenStudy (perl):

:)

OpenStudy (some.random.cool.kid):

so let me see if i understand this correctly ...

OpenStudy (some.random.cool.kid):

step 1: i divide 11.26 by 8 step 2:(1.05)^n final step: i fail the exam :P just kidding

OpenStudy (perl):

but you know what, lets keep it simple. the domain is just n >= 0

OpenStudy (perl):

number of days n is zero or positive, but not negative. that should be enough for reasonable domain

OpenStudy (some.random.cool.kid):

oh i think i got it plus i just looked up the domain again for better clarification. thanks :}

OpenStudy (some.random.cool.kid):

but before i close this is it ok if i ask a different question and then i tag you in? @perl

OpenStudy (perl):

ok

OpenStudy (some.random.cool.kid):

ok brb

OpenStudy (perl):

yes

OpenStudy (some.random.cool.kid):

i opened a new one

OpenStudy (some.random.cool.kid):

it starts with belinda